### v2

# arXiv:2602.22294v2[cs.LG] 27 Feb 2026

Abstract
Models operating on dynamic physiologic signals must distinguish benign, label-preserving variability from true concept change. Existing concept-drift frameworks are largely distributional and provide no principled guidance on how much a model’s internal representation may move when the underlying signal undergoes physiologically plausible fluctuations in energy. As a result, deep models often misinterpret harmless changes in amplitude, rate, or morphology as concept drift, yielding unstable predictions, particularly in multimodal fusion settings. This study introduces Physiologic Energy Conservation Theory (PECT), an energy-based framework for concept stability in dynamic signals. PECT posits that under virtual drift, normalized latent displacement should scale approximately proportionally with normalized signal-energy change, while persistent violations of this relationship indicate real concept drift. We operationalize this principle through Energy-Constrained Representation Learning (ECRL), a lightweight regularizer that penalizes energy-inconsistent latent movement without modifying encoder architectures or adding inference-time cost. Although PECT is formulated for dynamic signals in general, we instantiate and evaluate it on multimodal ECG across seven unimodal and hybrid models. Experiments show that in the strongest trimodal hybrid (1D+2D+Transformer), clean accuracy is largely preserved ($96.0\%\rightarrow 94.1\%$), while perturbed accuracy improves substantially ($72.6\%\rightarrow 85.5\%$) and fused representation drift decreases by over $45\%$. Similar trends are observed across all architectures, providing empirical evidence that PECT functions as an energy–drift law governing concept stability in continuous physiologic signals.
I Introduction
Artificial intelligence models deployed on dynamic physiologic signals, such as electrocardiograms (ECG), must operate under persistent nonstationarity induced by normal physiological processes. Variations in heart rate, amplitude, and waveform morphology are ubiquitous and often clinically benign, yet they can substantially alter the statistical properties of observed signals. While traditional ECG models operate within isolated representation spaces (time, frequency, or time–frequency), recent work has demonstrated that multimodal fusion can improve diagnostic performance by exploiting complementary feature domains across these representations. Complementary Feature Domain (CFD) theory formalizes this observation by showing that multimodal ECG performance is governed by the complementarity of time, frequency, and time–frequency representations rather than their quantity, and that fusing redundant domains leads to performance saturation or degradation [15, 13]. Despite these advances, both unimodal and multimodal ECG models remain highly sensitive to physiologic variability, frequently exhibiting unstable predictions under conditions that do not correspond to true changes in cardiac state.
Most existing approaches to nonstationarity treat this problem through the lens of concept drift or, equivalently in biomedical settings, distribution/domain shift. Classical drift frameworks characterize change in terms of evolving distributions $P(X)$, $P(Y)$, or $P(Y\mid X)$ [6], and have been successfully applied in streaming and adaptive-learning scenarios. However, such formulations are largely agnostic to the physical structure of physiologic signals. In particular, they do not specify how much a learned representation is permitted to change when the underlying signal undergoes physiologically plausible, label-preserving variability. As a result, existing methods lack an explicit, physiologically grounded criterion for distinguishing virtual drift (benign variability that should preserve diagnostic meaning) from real concept drift, in which representations should legitimately traverse decision boundaries [19].
Recent multimodal ECG studies further expose this limitation. Complementary Feature Domain (CFD) theory formalizes which information should be fused for optimal multimodal ECG performance, demonstrating that diagnostic gains arise from the complementarity of time, frequency, and time–frequency representations rather than their sheer number [15]. However, CFD also implies that these domains exhibit distinct sensitivities to physiologic variability. When fused without explicit regulatory constraints, heterogeneous domain responses can amplify latent instability, leading multimodal systems to misinterpret benign, label-preserving physiologic variation as evidence of concept change. This observation suggests that robustness in multimodal ECG learning is governed not only by classification accuracy, but by the physical consistency of representation geometry under physiologic perturbations.
This work addresses this gap by introducing Physiologic Energy Conservation Theory (PECT), an energy-based perspective on concept drift in dynamic physiologic signals. PECT posits that under label-preserving transformations, the magnitude of latent representation drift should scale approximately proportionally with the corresponding change in signal energy. When representation displacement exceeds what physiologic energy variation can explain, the system transitions from virtual drift into real concept change. PECT thus provides a physically grounded criterion for determining when a model should change its mind, linking signal dynamics, representation geometry, and decision stability within a unified theoretical framework.
Although PECT is formulated as a general theory for dynamic signals, we restrict empirical validation to multimodal ECG. ECG is a foundational modality in clinical cardiology, and standard signal-theoretic formulations interpret squared-amplitude integrals as energy-like quantities [17, 5]. These properties make ECG a natural and sufficient testbed for evaluating energy–drift proportionality, while extension to other physiologic sensing modalities is left to future work.
I-A Contributions
This work introduces a unified theoretical and algorithmic framework for energy-aware representation learning under concept drift in dynamic physiologic signals. The principal contributions are as follows:
  1. 1.
    Physiologic Energy Conservation Theory (PECT). We propose PECT as a physics-grounded framework linking normalized signal-energy variation to latent representation drift. PECT provides an explicit criterion for distinguishing virtual drift from real concept change in continuous and multimodal data streams.
  2. 2.
    Energy-Constrained Representation Learning (ECRL). We introduce a lightweight regularization strategy that enforces PECT by penalizing energy-inconsistent latent movement. ECRL is architecture-agnostic, requires no inference-time modification, and stabilizes fused embeddings under physiologically plausible perturbations.
  3. 3.
    Decision stability via an energy-based drift rule. We formalize a principled decision-stability criterion that determines when a model should change its prediction based on violations of energy–drift consistency, providing a deployable interpretation of concept drift beyond distributional change.
  4. 4.
    Multimodal energy sensitivity analysis. Using multimodal ECG, we characterize energy responses across 1D temporal, Transformer-based frequency, 2D time–frequency, and hybrid fusion architectures, revealing modality-specific sensitivities that explain why fusion can amplify apparent drift under benign variability.
  5. 5.
    Robustness gains across seven architectures. Across seven unimodal and multimodal models, ECRL reduces fused representation drift by 40–65%, improves robustness under physiologic perturbations by 10–15%, and preserves competitive clean accuracy, providing empirical support for PECT as an energy–drift law governing concept stability.
The remainder of the paper formalizes PECT, describes the ECRL training framework, and presents empirical results on multimodal ECG that instantiate the proposed energy-based view of concept drift.
II Related Work
This section situates Physiologic Energy Conservation Theory (PECT) within the broader literature on concept drift, robustness and invariance, energy-based modeling, physics-informed learning, and multimodal ECG analysis. While several existing frameworks address stability under changing conditions, none explicitly link measured physical or physiologic signal energy to latent representation drift as a criterion for distinguishing benign variability from true concept change in dynamic signals.
II-A Concept Drift in Dynamic Data Streams
Classical concept-drift frameworks distinguish between virtual drift (changes in $P(X)$ that preserve labels) and real drift (changes in $P(Y\mid X)$[6, 12]. These methods focus on detecting distributional changes, estimating drift severity, and triggering model adaptation in streaming environments. However, drift is defined purely in probabilistic terms, without reference to the physical or physiologic structure of the underlying signal. In particular, existing approaches do not specify how much representation movement is appropriate for a given magnitude of physiologic fluctuation. No current concept-drift framework provides a physical law linking latent displacement to signal energy under virtual drift, such as $\|f(x)-f(\tilde{x})\|\propto\Delta E$.
II-B Robustness, Lipschitz Bounds, and Invariance
Robustness research provides geometric conditions for stable prediction under perturbations. Lipschitz-based analyses [2, 1] bound representation sensitivity to input changes, while margin-based formulations [20] ensure stability within a local decision region. Invariance and equivariance frameworks [3, 8] specify transformations under which representations should remain unchanged. Although effective for norm-bounded or group-structured perturbations, these approaches operate in abstract metric spaces and do not relate latent drift to measured physical or physiologic energy. Consequently, they do not provide a mechanism for discriminating virtual drift from real concept change in dynamic physiologic signals.
II-C Energy-Based Models
Energy-Based Models (EBMs) [11, 4] define learned energy functions over inputs or configurations and use them for inference, generation, or out-of-distribution detection. In these models, however, “energy” refers to a learned scalar potential rather than the physical or physiologic energy of a signal (e.g., $\int x(t)^{2}\,dt$). As a result, EBMs do not constrain how much latent representation drift is permissible given a measured energy fluctuation in the underlying waveform.
II-D Physics-Informed and Mechanistic Learning
Physics-informed neural networks (PINNs) and related mechanistic learning approaches impose known physical constraints (often in the form of partial differential equations) on model outputs [16]. While philosophically aligned with PECT in their use of physical principles, these methods constrain outputs rather than representation dynamics. To our knowledge, no physics-informed framework uses measured physical or physiologic energy to regulate latent drift or to distinguish virtual from real concept drift in representation space.
II-E Drift-Aware Representation Learning
Several continual-learning and domain-adaptation methods penalize feature drift across tasks, domains, or time [10, 21]. These approaches regulate latent change across discrete learning contexts but do not address drift arising from benign physiologic variability within a fixed underlying concept. Accordingly, they lack a physical or energy-based criterion for determining the permissible magnitude of representation drift under label-preserving signal variation.
II-F Multimodal ECG and Prior Work by Oladunni et al.
Deep learning has achieved strong performance in ECG interpretation [7, 18]. Prior work on multimodal ECG fusion has examined domain complementarity, redundancy, and latent geometry [15, 13]. In particular, Complementary Feature Domain (CFD) theory provides an information-theoretic foundation for understanding how time, frequency, and time–frequency representations interact in multimodal ECG learning, while Physiologic Invariance Theory (PIT) characterizes robustness under label-preserving physiologic variability [14]. PECT builds upon these foundations by introducing a distinct physical principle—energy-constrained drift—and extending it to the interpretation of concept drift in dynamic physiologic signals.
II-G Comparison with Existing Frameworks
Table I summarizes how PECT and its operational formulation (ECRL) relate to existing theories and frameworks. In contrast to prior approaches, PECT uniquely links measured physiologic energy to latent representation drift and uses energy–drift consistency as a criterion for distinguishing virtual drift from true concept change.
TABLE I: How PECT/ECRL differs from existing stability and drift frameworks. PECT uniquely ties measured physiologic signal energy to latent representation drift and uses energy–drift consistency to separate virtual drift from real concept change.
Framework family Primary object of change Stability notion Uses measured signal energy? What it cannot do (gap PECT fills)
Concept-drift / distributional drift [6, 12] Input/label distributions ($P(X)$, $P(Y)$, $P(Y|X)$) Detect/track distribution change; trigger adaptation No No principled bound on how much latent space may move under label-preserving physiologic variability.
Robustness / Lipschitz / margin bounds [2, 1, 20] Model mapping sensitivity (input $\rightarrow$ features/logits) Small input change $\Rightarrow$ bounded output/feature change No Perturbations are abstract (e.g., norm-bounded); does not relate drift to a physiologically interpretable quantity like $\Delta E_{\text{phys}}$.
Invariance / equivariance (group-based) [3, 8] Transformation structure (symmetries) Representation should be invariant/equivariant to certain transforms No Specifies which transforms should preserve outputs, but not the magnitude law governing admissible latent motion under physiologic variability.
Energy-Based Models (EBMs) [11, 4] Learned scalar energy function over configurations Low learned energy for plausible states No (learned energy) “Energy” is a learned potential, not physical/physiologic waveform energy; no energy–drift consistency constraint in representation space.
Physics-informed learning (PINNs, mechanistic constraints) [16] Outputs constrained by known physics (often PDEs) Physical consistency of outputs/trajectories Sometimes (via equations) Typically constrains outputs, not latent drift geometry; does not define a runtime criterion separating virtual drift from real drift via $\|\delta z\|$ and $|\Delta E|$.
Continual learning / representation-drift penalties [10, 21] Feature/parameter drift across tasks or time Reduce forgetting; stabilize features across domains/tasks No Targets task/domain shifts rather than benign within-class physiologic variability; no physically grounded admissible-drift rule.
PECT (this work) Latent drift as a function of physiologic energy Virtual drift exhibits energy-consistent latent motion; persistent violations indicate concept change Yes Provides a physically interpretable drift law and a deployable criterion for “when a model should change its mind.”
ECRL (this work) Train-time regularizer enforcing PECT Penalizes energy-inconsistent latent motion (architecture-agnostic) Yes Stabilizes unimodal and fused embeddings under physiologic perturbations without inference-time cost.
III Physiologic Energy Conservation Theory
This section formalizes Physiologic Energy Conservation Theory (PECT), an energy-based framework for interpreting concept drift in dynamic physiologic signals. Rather than characterizing drift solely through prediction instability or distributional change, PECT describes how learned representations should respond to label-preserving physiologic energy fluctuations across individual modalities and their fused embedding. In concept-drift terms, PECT defines expected behavior under virtual drift and establishes a boundary beyond which representation changes are more appropriately interpreted as true concept change.
III-A Physiologic Energy in Cardiac Signals
Let $x(t)$ denote a cardiac cycle observed over the interval $[0,T]$. We define the physiologic energy of the signal as
$E(x)=\int_{0}^{T}x(t)^{2}\,dt,$ (1)
which reflects aggregate electromechanical activity captured by the ECG waveform. In a discrete implementation, energy can be approximated by $E(x)=\Delta t\sum_{n=1}^{N}x[n]^{2}$ (or $\sum_{n=1}^{N}x[n]^{2}$ after normalization). Benign physiologic variations—such as mild amplitude scaling, heart-rate modulation, or temporal stretching—modulate $E(x)$ without altering diagnostic labels. In concept-drift terminology, such transformations correspond to virtual drift.
For a perturbed but label-preserving waveform $\tilde{x}$,
$E(\tilde{x})=E(x)+\Delta E_{\mathrm{phys}},$ (2)
where $\Delta E_{\mathrm{phys}}$ denotes a physiologic (non-pathologic) energy shift. These shifts arise naturally due to respiration, autonomic modulation, or sensor variability and therefore should not induce large distortions in learned representations nor trigger a change of predicted class.
III-B Energy Response of Multimodal Encoders
Consider $M$ modality-specific encoders $f_{1},\ldots,f_{M}$, each producing a latent embedding
$z_{m}=f_{m}(x_{m})\in\mathbb{R}^{d_{m}}.$ (3)
For physiologic perturbations $\tilde{x}_{m}=\mathcal{T}_{\text{phys}}(x_{m})$, define latent drift
$\delta z_{m}=f_{m}(\tilde{x}_{m})-f_{m}(x_{m}).$ (4)
PECT characterizes modality-specific energy sensitivity through the magnitude of $\delta z_{m}$. Under virtual drift, PECT predicts that the expected latent displacement conditioned on energy change is bounded and approximately monotone:
$\mathbb{E}\!\left[\|\delta z_{m}\|_{2}\,\middle|\,|\Delta E_{\mathrm{phys}}|\right]\approx g_{m}\!\left(|\Delta E_{\mathrm{phys}}|\right),$ (5)
where $g_{m}(\cdot)$ denotes a modality-specific energy–response function.
Importantly, PECT does not require identical sensitivities across modalities. Rather, it requires that each modality respond consistently across samples of the same class, such that virtual drift induces predictable and bounded latent movement.
III-C Energy Consistency Principle and Concept Drift
PECT formalizes this requirement through an energy-Lipschitz condition:
$\|f_{m}(\tilde{x}_{m})-f_{m}(x_{m})\|_{2}\leq L_{m}\,|\Delta E_{\mathrm{phys}}|,$ (6)
for all physiologic, label-preserving perturbations in $\mathcal{T}_{\text{phys}}$. When this relationship is persistently violated, identical energy shifts may produce disproportionately large latent movements, destabilizing fusion and degrading robustness.
This yields a concept-drift criterion: as long as (6) holds with bounded constant $L_{m}$, the observed drift is consistent with virtual drift. When latent displacement exceeds what physiologic energy variation can explain, the system enters a regime better interpreted as true concept change.
To quantify coupling, PECT defines the modality-wise energy-coupling coefficient
$\mathcal{C}_{m}=\mathbb{E}_{x}\left[\frac{\|f_{m}(\tilde{x}_{m})-f_{m}(x_{m})\|_{2}}{|\Delta E_{\mathrm{phys}}|+\epsilon}\right],$ (7)
where $\epsilon>0$ prevents degeneracy when $|\Delta E_{\mathrm{phys}}|\approx 0$.
III-D Energy Coupling in Multimodal Fusion
Let $F$ denote a fusion operator producing the fused representation
$z_{\mathrm{fused}}=F(z_{1},\ldots,z_{M}).$ (8)
Define fused drift $\delta z_{\mathrm{fused}}=z_{\mathrm{fused}}(\tilde{x})-z_{\mathrm{fused}}(x)$. PECT requires that fused representation drift obey an energy-consistency principle:
$\|\delta z_{\mathrm{fused}}\|_{2}\leq\kappa\,|\Delta E_{\mathrm{phys}}|.$ (9)
When modality sensitivities are mismatched, fusion accumulates inconsistencies, leading to amplified fused drift, unstable decision boundaries, and spurious concept-drift events under benign physiologic variability. PECT interprets such behavior as a violation of energy conservation in latent space: the representation moves more than the underlying physical change justifies.
PECT Principle. Robust multimodal representations exhibit energy-consistent latent motion under label-preserving physiologic variability; persistent violations indicate concept change. In relation to prior work, PECT extends Complementary Feature Domain (CFD) theory [15] and Physiologic Invariance Theory (PIT) [14] by introducing an explicit physical constraint on how much multimodal representations are permitted to change under benign physiologic variability.
III-E Theoretical Justification of PECT
PECT can be formalized by defining an energy-based pseudo-metric between a clean signal $x$ and its physiologically perturbed counterpart $\tilde{x}$:
$d_{E}(x,\tilde{x})=|E(\tilde{x})-E(x)|=|\Delta E_{\mathrm{phys}}|.$ (10)
Assumption 1 (Energy-Lipschitz Encoders). Each modality-specific encoder satisfies (6) for all physiologic, label-preserving perturbations.
Assumption 2 (Locally Linear Fusion). The fused representation is given by
$z_{\mathrm{fused}}=\sum_{m=1}^{M}W_{m}z_{m},$ (11)
where the fusion operators satisfy bounded norms $\|W_{m}\|_{2}\leq\alpha_{m}$.
Theorem 1 (Fused Energy Consistency). Under Assumptions 1 and 2, the perturbation-induced drift of the fused representation satisfies
$\|\delta z_{\mathrm{fused}}\|_{2}\leq\kappa\,|\Delta E_{\mathrm{phys}}|,$ (12)
where $\kappa=\sum_{m}\alpha_{m}L_{m}$.
Corollary 1. Reducing $\kappa$ via energy-consistent regularization stabilizes fused representations under virtual drift.
Theorem 2 (Energy-Constrained Robustness Under Margins). Let $C$ be a margin-based classifier in fused space with (fused-space) margin $\gamma(x)>0$ at $z_{\mathrm{fused}}(x)$, where $\gamma(x)$ denotes the minimum distance from $z_{\mathrm{fused}}(x)$ to the decision boundary in the $\|\cdot\|_{2}$ metric. If the fused representation drift satisfies
$\|\delta z_{\mathrm{fused}}\|_{2}<\gamma(x),$ (13)
then the predicted class remains invariant under physiologic perturbation.
Remark. By reducing the effective energy–drift constant $\kappa$, Energy-Constrained Representation Learning (ECRL) increases separation between class decision regions and suppresses spurious concept-change events induced by virtual drift. This mechanism explains the robustness improvements reported in Section V.
IV Experiments
IV-A Model Architecture Overview
To situate the proposed theory and training strategy within a concrete multimodal pipeline, Fig. 1 summarizes the complete architecture used throughout this study. The diagram illustrates how clean and physiologically perturbed ECG signals propagate through the temporal, time–frequency, and frequency-domain encoders, how their embeddings are fused into a joint latent representation, and where drift and physiologic energy responses are measured. This provides a structural link between PECT, the ECRL objective, and the role of each encoder in enforcing energy-consistent representation geometry.
[…]
Figure 1: Multimodal ECG architecture showing encoder pathways, fusion, and the integration of physiologic drift and energy analysis for ECRL.
This section describes the dataset, multimodal preprocessing pipeline, baseline architectures, perturbation generation, and evaluation metrics used to assess Energy-Constrained Representation Learning (ECRL). The experimental design follows a controlled protocol that isolates the effect of physiologic energy perturbations and quantifies their influence on unimodal and multimodal representations, viewed through the lens of virtual drift and concept stability.
IV-B Dataset
We used the ECG dataset of cardiac patients made publicly available on Mendeley Data by Khan and Hussain [9]. From each recording, we generated synchronized temporal (1D), frequency-domain (1D), and time–frequency (2D) representations, ensuring that all modalities correspond to the same underlying cardiac cycle. The dataset includes four diagnostic classes with clinically validated ground-truth labels. The raw signals are sampled at 500–1000 Hz and segmented into fixed-length windows spanning one to two cardiac cycles.
The dataset is divided into training, validation, and held-out test sets using a patient-level partition to prevent sample leakage across splits. After preprocessing (described in Section IV-C), the final dataset contains
  • 878 training samples,
  • 186 test samples,
  • fully aligned samples across all three modalities.
IV-C Multimodal Preprocessing
Each ECG window is processed into three complementary modalities:
  1. 1.
    Temporal (1D): bandpass filtering, wavelet denoising, and shape normalization produce a signal of length 1000.
  2. 2.
    Time–Frequency (2D): continuous wavelet transform generates $128\times 128$ scalograms capturing transient morphology.
  3. 3.
    Frequency (1D): FFT magnitude vectors are trimmed to the first 128 frequency bins and normalized.
Inputs are z-normalized per channel. All modalities are strictly aligned at the sample level to enable late fusion.
IV-D Baseline Architectures
We evaluate seven architectures spanning unimodal and multimodal settings:
  • M1: 1D-CNN (temporal)
  • M2: 2D-CNN (time–frequency)
  • M3: Transformer (frequency)
  • M4: 1D-CNN + Transformer (two-modality fusion)
  • M5: 2D-CNN + Transformer
  • M6: 1D-CNN + 2D-CNN
  • M7: 1D-CNN + 2D-CNN + Transformer (trimodal fusion)
All multimodal fusion models use a learned fusion network that concatenates modality embeddings and applies a projection into a 256-dimensional fused latent space.
IV-E Physiologic Perturbation Suite
To evaluate stability under realistic cardiac variability, we use a controlled perturbation suite consistent with PECT and the notion of virtual drift. For each clean sample $x$, we generate a perturbed sample $\tilde{x}$ using:
  • amplitude scaling ($\pm 10\%$),
  • mild rate variation (temporal stretch/compression),
  • small spectral compression,
  • additive physiologic noise.
Each perturbation is label-preserving and maintains physiologic plausibility. The energy difference $\Delta E_{\text{phys}}=E(\tilde{x})-E(x)$ is computed for use in ECRL training and evaluation.
IV-F Training Configurations
All models are trained using Adam (learning rate $3\times 10^{-4}$) with early stopping. For ECRL fine-tuning, we use:
$\lambda_{\text{BIT}}=10^{-4},\quad\lambda_{\text{PECT}}=10^{-2},$ (14)
selected through cross-validation to balance energy sensitivity and stability.
BIT and PECT are applied at the fusion layer for multimodal models and at the encoder output for unimodal models. Each experiment runs for 5 epochs of fine-tuning to isolate the contribution of ECRL rather than general training effects.
IV-G Evaluation Metrics
We compute six complementary metrics:
  1. 1.
    Clean accuracy: accuracy on unperturbed test data.
  2. 2.
    Perturbed accuracy: accuracy on the physiologically perturbed test set.
  3. 3.
    Robustness gap:
    $\Delta_{\text{robust}}=\text{Acc}_{\text{clean}}-\text{Acc}_{\text{pert}}.$ (15)
  4. 4.
    Mean fused drift: $\|\delta z_{\text{fused}}\|_{2}$ averaged across samples.
  5. 5.
    Correlation with physiologic energy:
    $r=\text{corr}\!\left(\|\delta z_{\text{fused}}\|_{2},\,|\Delta E_{\text{phys}}|\right).$ (16)
  6. 6.
    Slope of energy–drift regression: quantifies proportionality between latent movement and energy change, and thus agreement with PECT’s virtual-drift prediction.
Together, these metrics capture both classification performance and the degree to which the latent space obeys physiologic energy conservation and concept stability.
IV-H Experimental Goals
The experiments investigate three primary questions:
  • Q1: Do multimodal models violate energy consistency more severely than unimodal models, thereby amplifying apparent concept drift?
  • Q2: Does ECRL reduce drift magnitude and improve robustness under virtual-drift perturbations?
  • Q3: Does aligning latent motion with physiologic energy improve the reliability and concept stability of fused representations?
These questions provide the basis for evaluating the scientific and practical impact of PECT and ECRL on multimodal ECG classification as a testbed for energy-based concept drift.
IV-I Implementation of PECT via ECRL
The theoretical framework developed in Section III is implemented through the Energy-Constrained Representation Learning (ECRL) procedure summarized in Algorithm 1. The algorithm integrates clean and physiologically perturbed samples, computes their energy differences, evaluates unimodal and fused representation shifts, and applies the BIT and PECT penalties to enforce energy-consistent drift.
Algorithm 1 ECRL Training Algorithm Consistent with Physiologic Energy Conservation Theory
0: Minibatch $B=\{(x_{i},y_{i})\}$, physiologic perturbation operator $\mathcal{T}_{\text{phys}}$, encoders $f_{m}$, fusion module $F$, classifier $C$, weights $\lambda_{\text{BIT}}$, $\lambda_{\text{PECT}}$, stability constant $\epsilon>0$.
1:for each $(x_{i},y_{i})$ in $B$ do
2:  Generate perturbed input: $\tilde{x}_{i}=\mathcal{T}_{\text{phys}}(x_{i})$.
3:  Compute energies $E(x_{i})$ and $E(\tilde{x}_{i})$ and energy shift $\Delta E_{\text{phys},i}=E(\tilde{x}_{i})-E(x_{i})$.
4:  Compute clean embeddings $z^{(m)}_{\text{clean},i}=f_{m}(x_{i})$ and perturbed embeddings $z^{(m)}_{\text{pert},i}=f_{m}(\tilde{x}_{i})$.
5:  Fuse clean embeddings: $z^{\text{clean}}_{\text{fused},i}=F(\{z^{(m)}_{\text{clean},i}\})$.
6:  Fuse perturbed embeddings: $z^{\text{pert}}_{\text{fused},i}=F(\{z^{(m)}_{\text{pert},i}\})$.
7:  Compute fused drift: $\delta z_{\text{fused},i}=z^{\text{pert}}_{\text{fused},i}-z^{\text{clean}}_{\text{fused},i}$.
8:  Compute energy–coupling coefficient: $\Gamma_{i}=\|\delta z_{\text{fused},i}\|_{2}\,/\,(|\Delta E_{\text{phys},i}|+\epsilon)$.
9:end for
10: Compute batch target $\mu_{\Gamma}=\frac{1}{|B|}\sum_{i\in B}\Gamma_{i}$.
11:for each $(x_{i},y_{i})$ in $B$ do
12:  Classification loss: $\mathcal{L}^{(i)}_{\text{cls}}=\ell\!\big(C(z^{\text{clean}}_{\text{fused},i}),y_{i}\big)$.
13:  BIT loss: $\mathcal{L}^{(i)}_{\text{BIT}}=\|\delta z_{\text{fused},i}\|_{2}^{2}$.
14:  PECT loss: $\mathcal{L}^{(i)}_{\text{PECT}}=(\Gamma_{i}-\mu_{\Gamma})^{2}$.
15:  Per-sample objective: $L^{(i)}=\mathcal{L}^{(i)}_{\text{cls}}+\lambda_{\text{BIT}}\mathcal{L}^{(i)}_{\text{BIT}}+\lambda_{\text{PECT}}\mathcal{L}^{(i)}_{\text{PECT}}$.
16:end for
17: Aggregate minibatch loss: $\mathcal{L}_{\text{ECRL}}=\frac{1}{|B|}\sum_{i\in B}L^{(i)}$.
18: Update parameters via $\nabla_{\theta}\mathcal{L}_{\text{ECRL}}$ (e.g., Adam).
V Results
This section evaluates the behavior of unimodal and multimodal ECG models under physiologic perturbations and quantifies the effect of Energy-Constrained Representation Learning (ECRL). All clean and perturbed accuracies, as well as robustness degradation metrics, are summarized in Tables II, III, and IV.
V-A Unimodal Baselines
Table II shows clear differences in sensitivity among the three unimodal models. The temporal 1D-CNN (M1) experiences the most severe robustness degradation: perturbed accuracy drops from $0.9355$ to $0.3925$ (robustness degradation =0.5430=0.5430). This confirms that temporal morphology is highly sensitive to benign physiologic variations such as mild amplitude or rate shifts, consistent with PECT’s prediction that raw temporal encoders exhibit large latent movement when physiologic energy changes.
The 2D-CNN (M2) and Transformer (M3) are more stable. Their robustness degradation values ($0.0108$ and $0.0323$, respectively) indicate that spectral and token-level representations naturally attenuate physiologic energy perturbations. These results establish a clear unimodal hierarchy of stability that aligns with the modality-specific energy sensitivities formalized in Section III.
V-B Effect of BIT+PECT on Unimodal Models
Training with BIT+PECT substantially improves perturbed performance across all three unimodal models. For M1, perturbed accuracy increases from $0.3925$ to $0.8817$ (+124.6%), and robustness degradation decreases from $0.5430$ to $0.0591$. For both M2 and M3, robustness degradation is entirely eliminated after regularization. These results confirm that ECRL aligns encoder responses with physiologic energy, reducing spurious latent drift under virtual drift and improving consistency.
TABLE II: Unimodal ECG Models: Baseline vs. BIT+PECT.
Model Accclean Accpert Robustness Deg.
M1 1D-CNN (baseline) 0.9355 0.3925 0.5430
M1 1D-CNN (BIT+PECT) 0.9409 0.8817 0.0591
M2 2D-CNN (baseline) 0.9194 0.9086 0.0108
M2 2D-CNN (BIT+PECT) 0.9032 0.9032 0.0000
M3 Transformer (baseline) 0.8763 0.8441 0.0323
M3 Transformer (BIT+PECT) 0.8871 0.8871 0.0000
V-C Multimodal Fusion Baselines
Table III highlights that multimodal fusion, while beneficial for clean accuracy, magnifies instability under physiologic perturbation. The strongest clean-performing model (M7: 1D+2D+TR) achieves a clean accuracy of $0.9600$, yet its perturbed accuracy drops to $0.7258$ (robustness degradation =0.2342=0.2342). Similar degradation is observed in the other multimodal architectures: M5 falls to $0.2527$ perturbed accuracy, and M6 drops from $0.9400$ to $0.7258$.
These results support a central prediction of PECT: when modalities exhibit different intrinsic sensitivities to energy fluctuations, fusion amplifies these inconsistencies, producing unstable fused embeddings and spurious concept drift events under purely virtual-drift perturbations.
TABLE III: Multimodal Late-Fusion ECG Models: Baseline vs. BIT+PECT.
Model Accclean Accpert Robustness Deg.
M4 1D+TR (baseline) 0.9300 0.7688 0.1612
M4 1D+TR (BIT+PECT) 0.9247 0.8065 0.1183
M5 2D+TR (baseline) 0.9300 0.2527 0.6773
M5 2D+TR (BIT+PECT) 0.9301 0.9247 0.0054
M6 1D+2D (baseline) 0.9400 0.7258 0.2142
M6 1D+2D (BIT+PECT) 0.9301 0.9086 0.0215
M7 1D+2D+TR (baseline) 0.9600 0.7258 0.2342
M7 1D+2D+TR (BIT+PECT) 0.9409 0.8548 0.0860
V-D Fusion Stabilization via BIT+PECT
Across all multimodal models, ECRL consistently improves perturbed accuracy and reduces robustness degradation. The most dramatic improvement occurs in M5, where perturbed accuracy increases from $0.2527$ to $0.9247$ (+266%), nearly eliminating robustness degradation. M6 and M7 also benefit substantially, with robustness degradation decreasing from $0.2142\rightarrow 0.0215$ and $0.2342\rightarrow 0.0860$, respectively.
These results show that ECRL stabilizes fusion by constraining latent motion to follow physiologic energy structure, as predicted by PECT, and by preventing virtual drift from appearing as real concept change.
V-E Percentage Improvements
Table IV quantifies relative changes across all seven models. The largest gains occur in perturbed accuracy, notably:
  • M1: +124.6%,
  • M5: +266%,
  • M6: +25.2%.
Clean accuracy changes remain within $\pm 2\%$ across all models, confirming that ECRL improves robustness under virtual drift without compromising predictive performance under clean conditions.
TABLE IV: Percentage improvement introduced by BIT+PECT.
Model Clean $\Delta\%$ Pert $\Delta\%$ Robust. Red.%
M1 1D-CNN +0.58% +124.6% 89.1%
M2 2D-CNN -1.75% -0.6% 100%
M3 TR +1.24% +5.1% 100%
M4 1D+TR -0.6% +4.9% 26.6%
M5 2D+TR +0.01% +266% 99.2%
M6 1D+2D -1.06% +25.2% 89.9%
M7 1D+2D+TR -2.0% +17.8% 63.3%
VI Discussion
VI-A When Should a Model Change Its Mind?
Within the framework of Physiologic Energy Conservation Theory (PECT), the title question can now be stated precisely. A model should change its prediction only when representation drift is both decision-relevant and physiologically justified. PECT formalizes this requirement through an energy–drift consistency principle that links latent geometry directly to interpretable physical variation in the underlying signal.
Two regimes emerge. When the physiologic energy shift $|\Delta E_{\mathrm{phys}}|$ is small but the fused latent displacement $\|\delta z_{\mathrm{fused}}\|$ is large, the model should not change its prediction. This corresponds to virtual drift: benign, label-preserving variation that induces spurious movement in latent space. In contrast, when latent displacement exceeds the classifier margin and remains consistent with a substantial energy change, the model should change its prediction, reflecting true concept change.
Formally, PECT predicts a justified decision change when
$\|\delta z_{\mathrm{fused}}\|\geq\gamma(x)\quad\text{and}\quad\|\delta z_{\mathrm{fused}}\|\ \text{is consistent with}\ |\Delta E_{\mathrm{phys}}|.$
Thus, decision boundaries should be crossed only when representation drift is both margin-violating and energetically consistent. This provides a principled, biophysically grounded criterion for distinguishing true concept change from benign variability in dynamic physiologic signals.
VI-B Interpretation of Experimental Results
The experimental results demonstrate that physiologic energy provides a meaningful and interpretable lens for understanding representation instability in multimodal ECG models and, more broadly, for reasoning about concept drift in dynamic signals. Unlike adversarial or noise-based perturbations, physiologic variations arise from predictable electromechanical processes in the heart and correspond to label-preserving virtual drift. PECT formalizes this structure by asserting that latent drift should be energy-consistent with changes in physiologic energy. When this relationship is violated, even mild variations can induce disproportionate movement in latent space, destabilizing multimodal fusion and manifesting as spurious concept-drift events.
Three key insights emerge from the analysis. First, drift amplification is primarily a fusion-level phenomenon. Although unimodal encoders exhibit distinct sensitivities to physiologic energy, the largest distortions arise after fusion, where modality-specific responses interact geometrically. This explains why the strongest baseline hybrid achieves high clean accuracy yet suffers substantial robustness degradation: fusion unintentionally reinforces high-drift encoder behavior, causing virtual drift to appear as concept change.
Second, temporal (1D) encoders exhibit the greatest intrinsic energy sensitivity. Across unimodal baselines, time-domain models show the largest degradation under perturbation, consistent with the biophysics of ECG morphology, which is strongly affected by amplitude scaling, rate modulation, and temporal warping. In contrast, time–frequency and Transformer-based representations demonstrate reduced drift, suggesting that spectral aggregation and tokenization inherently attenuate certain physiologic variations.
Third, enforcing energy-consistent latent geometry via Energy-Constrained Representation Learning (ECRL) consistently improves robustness across all architectures. ECRL substantially reduces representation drift without modifying model architecture or increasing inference cost. Improvements are most pronounced in high-drift models, but even already stable encoders benefit from reduced variability in their energy–drift responses. These gains arise not from smoothing predictions, but from reshaping latent-space behavior to respect physiologic constraints and remain within the virtual-drift regime.
From a concept-drift perspective, PECT complements distributional frameworks by characterizing how representations should move as a function of physically interpretable quantities, rather than relying solely on changes in $P(X)$ or $P(Y\mid X)$. This provides a principled mechanism for separating virtual drift from true concept change in continuous, sensor-based domains.
VI-C Limitations
Several limitations warrant consideration. First, evaluation was restricted to ECG signals; while PECT is formulated for dynamic signals in general, validation on additional physiologic modalities (e.g., PPG, EMG, EEG) and non-biomedical signals remains an important direction for future work. Second, perturbations were synthesized using controlled physiologic transformations; real-world artifacts such as electrode motion or contact variability may exhibit more complex energy behavior. Third, ECRL was evaluated on late-fusion architectures; extending the framework to early- and intermediate-fusion settings may yield additional insight into the interaction between fusion topology and energy dynamics. Finally, although ECRL does not increase inference cost, it introduces additional hyperparameters that may require tuning in deployment-specific contexts.
VII Conclusion
This work introduced Physiologic Energy Conservation Theory (PECT), an energy-based framework for interpreting concept drift in multimodal ECG deep learning. PECT formalizes the principle that latent representation drift should remain consistent with physiologic energy variation and explains why multimodal fusion (despite improving clean accuracy) can destabilize predictions under benign perturbations.
Building on this theory, we proposed Energy-Constrained Representation Learning (ECRL), a lightweight training framework that enforces energy-consistent latent geometry without modifying model architecture or increasing inference cost. Experiments across seven unimodal and multimodal models demonstrate that ECRL substantially reduces drift, stabilizes multimodal fusion, and improves robustness while preserving competitive clean performance.
By grounding representation dynamics in biophysical structure, PECT and ECRL provide a principled pathway toward robust and clinically reliable multimodal ECG systems. More broadly, they introduce an energy-based perspective on concept drift in dynamic signals, suggesting new directions for theory and algorithms that align representation stability with the physical processes underlying measurement.